OpenStudy (anonymous):

A 30.0 μF capacitor is initially charged with 15.0 mC. It is then discharged through a 200.0 Ω resistor. What is the maximum current during the discharge?

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

i havent delved into electrical stuff to know if im working it right or not. id have to google for an idea to run with

OpenStudy (amistre64):

something like this http://hyperphysics.phy-astr.gsu.edu/hbase/electric/capdis.html

OpenStudy (amistre64):

i = Q' Q = C Vb e^(-t/(RC)) Q' = -Vb/R e^(-t/(RC)) i beleive R is resistance, Vb is initial voltage charge, and C might be the capacitor info

OpenStudy (anonymous):

\[I _{0}=\frac{ Q _{0} }{ C*R }=2.5A\]

OpenStudy (anonymous):

i found answer thx